Reports 1-1 of 1 Clear search Modify search
AdV-SGD (FDS commissioning)
garaventa, polini, sorrentino - 12:36 Thursday 13 January 2022 (54467) Print this report
Stray light active loop up to FC IM: compensation of 500Hz dither line with LO actuator

Yesterday afternoon/evening we worked on the active control loop of scattered light going up to the input mirror of the filter cavity. We managed to compensate the 500Hz dither line put on M4_Z.

1. Visibility check and optimization

At the beginning of the shift the visibility on the HD between BAB and LO was ~36%.

Power check on PD1 HD:

  • LO: 0.771 V;
  • BAB: 0.438 V.

We moved M13 on SQB1 and M4/M6 on EQB1 and we reached ~75% (normalized).

2. Stray light loop closed without line compensation

Yesterday afternoon/evening the bump was smaller than last time (entry 54442).

We set the line amplitude to 15mV.

We managed to close the loop correcting the bump, see Fig. 1.

We tried to implement a boost in the correction filter:

  • GPS 16h24m00 UTC: loop open with 15mV line;
  • GPS 16h30m00 UTC: NO BOOST, loop closed with 15mV line, GAIN = 400;
  • GPS 16h36m48 UTC: NO BOOST, loop closed with 15mV line, GAIN = 500;
  • GPS 16h41m00 UTC: BOOST (pole 0.05 0, zero 5 0), loop closed with 15mV line, GAIN = 1800;
  • GPS 16h50m00 UTC: BOOST (pole 0.01 0, zero 10 0), loop closed with 15mV line, GAIN = 2500;
  • GPS 16h51m40 UTC: NO BOOST, loop closed with 15mV line, GAIN = 500.

In Fig. 2, magenta = with boost GAIN = 2500 (pole 0.01 0, zero 10 0) and blue without boost: the effect of the boost at low frequency is negligible.

3. Compensation of the 500Hz line engaging CC fast and CC coarse loops

We are in BAB mode.

In the entry 54344, we compensated the 500Hz line with the delay line, with a modulation on LO_M2 of 36mV amplitude and phi0 = 4.45. Since we were on the delay line, only the CC fast was needed and it was closed with a GAIN = 5.

Since yesterday we operated the system up to the input mirror of the filter cavity, we needed both the CC fast and CC coarse loops.

It was not possible to close the CC coarse on M4_Z so we sent the actuation on M6_Z.

We closed:

  • CC fast with GAIN = 10;
  • CC coarse with GAIN = -1.

Still sending a line of 15mV on M4_Z, we added a sine driving noise on the CC fast actuator (LO_M2) with frequency = 500 Hz and amplitude = 108mV (3x36mV, the optimal amplitude with a line of 5mV on M4_Z).

NB. In this configuration, changing the CC loop phase is worsing the balancing of the PDs of HD. This is expressed in a lowering of the contrast on HD.

We noticed that the PDs of HD were saturating (power > 2 V) so we moved the wave plate of EQB1 FI by 8000 steps.

After moving it we had:

  • BAB: 0.32V on PD2 HD (UTC 21h07m)

In order to compensate the line we optimized the amplitude and the phase of the sine sent to LO_M2.

We started with optimizing the parameters for 5mV on M4_Z and then we increased the line up to 15mV. We changed the phi0 on the driving noise sent to the fast actuator (phi of sinewave).

  • M4_Z line = 5mV, LO_M2 line = 36 mV -> phi0 = -1.8, see Fig. 3: the blue line is with a random phase while the magenta one is with the optimal phase.
  • M4_Z line = 15mV, LO_M2 line = 100 mV -> phi0 = -1.7, see Fig. 4: the magenta curve is with a quasi optimized phase (-1.8 rad) and an amplitude of 108 mV; the blue curve corresponds to the optimal phase (-1.7 rad) and the optimal amplitude (100mV).

In Fig. 7, we show the signals of the CC loops.

4. Stray light loop closed with compensation line on LO_M2

We turned off the BAB (so turned off also CC loops) and we engaged the stray light active loop with the compensation line on LO_M2 ON.

We closed it with M4_Z line at 15mV (compensation line at 100mV). In Fig. 5 we see that the loop is correcting the bump.

We closed it with M4_Z line at 5mV (compensation line at 36mV). In Fig. 6 we see that the loop is correcting the bump.

We conclude that the loop is working also when the compensation line on LO_M2 is ON.

Images attached to this report
Search Help